Solve the system by graphing. See Example 4. (Objective 3)
x = 5 − 2y
3x + 6y = 15
1. State whether the system is inconsistent or if the equations are dependent.
inconsistent
dependent
Correct: Your answer is correct.
2. Write the solution set. (Simplify your answer completely. If there is no solution, enter NO SOLUTION. If the system is dependent, enter a general solution in terms of x or y.)

Found 2 solutions by ewatrrr, ikleyn:
Answer by ewatrrr(24785)


Hi,
Solve the system by graphing. 

x = 5 − 2y    0r  y = -x/2 + 5/2   Green Line (slants to left - negative slope)

3x + 6y = 15  0r  y = x/2 + 15/6

  Ordered Pair P(0, 2.5) is the Solution for this system of Equations
 

A consistent system has at least one solution.
An inconsistent system has no solutions.  
An independent system has exactly one solution. ****
A dependent system has infinitely many solutions.

Wish You the Best in your Studies.


Answer by ikleyn(52937)

.

            The plots presented in the post by @ewatrrr are     IRRELEVANT     to the given problem.

            Her answer about the set of the solutions to the given problem is     INCORRECT.

            At this forum,     this woman systematically mislead visitors,     providing     INCORRECT     solutions.


--------------

Dear @ewatrrr.


PLEASE revise your solution.


I just tired to correct after you.